A canoeist paddled 6 miles upstream in 6 hours and returned to his starting point downstream in 45 minutes. What was the speed of the current?
Answer by mananth(16949)

A canoeist paddled 6 miles upstream in 6 hours and returned to his starting point downstream in 45 minutes. What was the speed of the current?
distance = same both ways
let speed of canoe be x mph
current speed = y mph
..
speed = distance/time
against current x-y = 6/6
with current x+y = 6/(3/4)===>6*4/3===>8
x-y =1
x+y=8
Add the two equations
2x=9
x= 4.5 mph speed of boat
speed of current = 8-4.5 = 3.5 mph